# Approximate independence of permutation mixtures

### v2

@@ 4.1 Divergence as a matrix permanent
Compared with Lemma 4.3, the result of Lemma 4.5 has an additional $O(\sqrt{\ell})$ factor, due to the use of Theorem 3.3 in the complex case. However, such an individual bound is crucial to deal with weighted sums $\sum_{\ell=0}^{n}w_{\ell}S_{\ell}$ in applications such as Theorem 1.6. Again, the proof of Lemma 4.5 relies on the use of complex normal random variables. Recalling the notation $A=UDU^{\top}$ from Section 4.1, let $\widetilde{U}\in\mathbb{R}^{n\times(n-1)}$ be the matrix $U$ with its first column (which is $\frac{\bf 1}{\sqrt{n}}$ by Lemma 4.2) removed, and $\widetilde{D}=\mathrm{diag}(\lambda_{2},\ldots,\lambda_{n})$ with the first eigenvalue $\lambda_{1}=1$ removed. The following lemma establishes an alternative expression of $S_{\ell}$.
@@
The proof of Lemma 4.6 again follows from the Isserlis’ theorem and several permutation identities in Lemma A.1, and is deferred to Section C.8. Now we present the proof of Lemma 4.5.
Proof of Lemma 4.5.
Since the matrix $U$ is orthogonal with the first column proportional to ${\bf 1}$, all columns of $\widetilde{U}$ are orthogonal to ${\bf 1}$. Consequently, the vector $\widetilde{P}z=\widetilde{U}\widetilde{D}^{1/2}z$ sums to zero, so (24) in Theorem 3.3 gives
@@
Next we use the individual upper bounds in Lemma 4.5 to prove Theorem 1.6. Similar to (26) and (27), let the matrix $A\in\mathbb{R}^{(n-1)\times(n-1)}$ be constructed based on $P_{2},\ldots,P_{n}\in{\mathcal{P}}$, and $S_{\ell}$ be the $-\ell$ homegeneous polynomial in its eigenvalues $(\lambda_{2},\ldots,\lambda_{n-1})$. Applying a similar likelihood ratio computation to Lemma 4.1 leads to the following upper bound.
Lemma 4.7.
Under the above notations, the following inequality holds:
${\sf TV}(\mathbb{P}_{n},\mathbb{Q}_{n})^{2}\leq\frac{1}{4}\int\frac{(\mathrm{d}\mathbb{P}_{n}-\mathrm{d}\mathbb{Q}_{n})^{2}}{\mathrm{d}\overline{P}^{\otimes n}}\leq\frac{{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})}{n}\sum_{\ell=1}^{n}\ell S_{\ell-1}.$
The proof of Lemma 4.7 is postponed to the appendix, and we proceed to prove Theorem 1.6. By Lemma 4.5, we have
where (a) is due to $\prod_{i=2}^{n-1}(1+\ell_{i})\geq 1+\sum_{i=2}^{n-1}\ell_{i}=\ell$, (b) uses $22\sqrt{\ell+1}\leq\ell+$ 2 for all $\ell\geq 0$, and (c) follows from the identity
By the same argument used to deal with the final product in the proof of Lemma 4.3, Lemma 4.7 leads to
${\sf TV}(\mathbb{P}_{n},\mathbb{Q}_{n})^{2}\leq\frac{3{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})}{n}(e\mathsf{\Delta}_{H^{2}}({\mathcal{P}}))^{3{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})},$
which establishes the first claim of Theorem 1.6. The second claim of Theorem 1.6 simply follows from the Cauchy–Schwarz inequality.
@@ 5.1 Tightness of upper bounds
In this section we discuss the tightness of the upper bounds in Theorem 1.2. Recall that the upper bound (4) is quadratic in ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})$ when ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})$ is small, and the upper bound (5) or (6) is a power function of $\mathsf{\Delta}_{H^{2}}({\mathcal{P}})$ when ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})$ is large. The next lemma shows that this is essentially the right behavior of $\chi^{2}(\mathbb{P}_{n}\|\mathbb{Q}_{n})$ for every family ${\mathcal{P}}$.
@@
In the setting of Theorem 1.2, the following lower bound holds:
@@
Combined with Theorem 1.2, Lemma 5.1 shows that when the capacity of the family ${\mathcal{P}}$ is small, we have ${\mathsf{D}}_{H^{2}}({\mathcal{P}})^{2}\lesssim\sup_{n,P_{1},\dots,P_{n}}\chi^{2}(\mathbb{P}_{n}\|\mathbb{Q}_{n})\lesssim{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})^{2}$, which is often tight as ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})\leq{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})\approx{\mathsf{D}}_{H^{2}}({\mathcal{P}})$ in view of the fact that most $f$-divergences are locally $\chi^{2}$-like [PW24, Chapter 7.10]. At the other extreme, when the family ${\mathcal{P}}$ is rich, we have $\mathsf{\Delta}_{H^{2}}({\mathcal{P}})^{\Omega(1)}-1\leq\sup_{n,P_{1},\dots,P_{n}}\chi^{2}(\mathbb{P}_{n}\|\mathbb{Q}_{n})\leq(e\mathsf{\Delta}_{H^{2}}({\mathcal{P}}))^{{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})}-1$. Therefore, up to a gap in the exponent, the maximum $H^{2}$ singularity $\mathsf{\Delta}_{H^{2}}({\mathcal{P}})$ plays a central role when it is large. The proof of Lemma 5.1 computes the variance of a linear test function $\frac{1}{n}\sum_{i=1}^{n}f(X_{i})$ under $\mathbb{P}_{n}$ and $\mathbb{Q}_{n}$.
The next natural question is on the tightness of the exponent ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})$ in the upper bound (5). Using the permanent representation of the $\chi^{2}$ divergence in Lemma 4.1, we can equivalently ask whether the following bound in Section 4 is tight:
@@
By Lemma 5.2, the exponent ${\mathsf{C}}$ in both the permanent upper bound (29) and the $\chi^{2}$ divergence upper bound (5) in Theorem 1.2 is tight in the worst case, up to multiplicative constants. However, as opposed to Lemma 5.1, the statement of Lemma 5.2 is not pointwise in ${\mathcal{P}}$; in other words, this does not mean that the upper bounds of Theorem 1.2 are always tight for a specific class ${\mathcal{P}}$. For example, for the Gaussian family (cf. point 1 of Corollary 1.3) with $\mu>1$, Theorem 1.2 and Lemma 5.1 only imply that $\exp(\Omega(\mu^{2}))\leq\sup_{n,P_{1},\dots,P_{n}}\chi^{2}(\mathbb{P}_{n}\|\mathbb{Q}_{n})\leq\exp(O(\mu^{3}))$, still exhibiting a gap on the exponent. An intuitive reason is that, for specific ${\mathcal{P}}$, further properties of the matrix $A$ (besides the trace and spectral gap) could potentially be exploited to lead to better results.
Although we leave the exploitation of further properties of $A$ as an open question, we mention one possible strategy here. Note that the upper bound in Section 4.2 claims that
$\chi^{2}(\mathbb{P}_{n}\|\mathbb{Q}_{n})+1\leq\prod_{i=2}^{n}\frac{1}{1-\lambda_{i}}=\left(n\sum_{T}\prod_{(i,j)\in T}A_{ij}\right)^{-1},$
where we sum over the edge set $T$ of all $n^{n-2}$ spanning trees of the complete graph over $[n]$, and the equality is due to the weighted matrix tree theorem [Kir47] applied to the graph Laplacian $I-A$. We conjecture that lower bounding this spanning tree polynomial of $A$ could lead to an improved upper bound for $\chi^{2}(\mathbb{P}_{n}\|\mathbb{Q}_{n})$ in the Gaussian family.
5.2 Application to a Gaussian sequence model
The contiguity relation $\{\mathbb{P}_{n}\}\triangleleft\{\mathbb{Q}_{n}\}$ in Theorem 1.2, or more specifically the mutual information relation (1) in Corollary 1.4, can be applied in statistical problems. We review a classical example of a Gaussian sequence model as follows. Let the observation vector be $X\sim{\mathcal{N}}(\theta,I_{n})$, with an unknown mean vector $\theta\in\mathbb{R}^{n}$ lying in the following $\ell^{p}$ ball $\Theta_{p}(R)$:
It is a classical result [DJ94] that for $q\geq p\vee 1$, the minimax risk is $r_{p,q}^{\star}(n,R)=(1+o(1))\mu^{q-p}R^{p}$ if $R=o(1)$ and $\frac{nR^{p}}{\mu^{p}}=\omega(1)$, where $\mu=\sqrt{2\log(1/R^{p})}$ is the threshold used in the soft-thresholding estimator. The quantity $s:=\frac{nR^{p}}{\mu^{p}}$ denotes the average sparsity level of $\theta$, and the extra condition $s=\omega(1)$ is used in the arguments of the minimax lower bound: when lower bounding the Bayes risk under the i.i.d. two-point prior [^3] In [DJ94], a three-point prior $(1-\alpha)\delta_{0}+\frac{\alpha}{2}(\delta_{-\mu}+\delta_{\mu})$ was used. However, their argument gives the same lower bound under the two-point prior. $(1-\alpha)\delta_{0}+\alpha\delta_{\mu}$ with $\alpha=(1+o(1))\frac{s}{n}$, the condition $s=\omega(1)$ ensures that the i.i.d. prior concentrates on $\Theta_{p}(R)$ with high probability. This condition was later removed in [Zha12], by lower bounding the Bayes risk under a permutation prior $\theta=(v_{\pi(1)},\dots,v_{\pi(n)})$ with $\pi\sim\text{\rm Unif}(S_{n})$ when $s=O(1)$, with a vector $v=(\mu,\dots,\mu,0,\dots,0)$ up to a rounding error under possibly non-integer $s$. The major benefit of the permutation prior is that $\theta\in\Theta_{p}(R)$ almost surely. However, based on the technical tools in [Zha12, Prop. 2.3], the current lower bound for permutation priors only works for a low sparsity level $s=O(\sqrt{n})$.
It is then a natural question to ask if the minimax risk $r_{p,q}^{\star}(n,R)=(1+o(1))\mu^{q-p}R^{p}$ can also be established using the above permutation mixture for all $s=\omega(1)$. It turns out that a combination of Corollary 1.4 and the Gaussian I-MMSE formula [GSV05] directly provides an affirmative answer for the squared loss $q=2$.
@@ 5.2 Application to a Gaussian sequence model
Let $q=2\geq p$ and $R=o(1)$. For $\mu=\sqrt{2\log(1/R^{p})}$ and $s=\frac{nR^{p}}{\mu^{p}}=\omega(1)$, the Bayes $\ell^{2}$ risk under the permutation prior $\theta=(v_{\pi(1)},\dots,v_{\pi(n)})$ with $\pi\sim\text{\rm Unif}(S_{n})$ and
is lower bounded by $(1-o(1))(1-c(\varepsilon))\mu^{2-p}R^{p}$, with $c(\varepsilon)\to 0$ as $\varepsilon\to 0^{+}$.
Choosing $\varepsilon\to 0^{+}$, Lemma 5.3 recovers the lower bound of $r_{p,2}^{\star}(n,R)$ via a permutation prior.
5.3 Empirical Bayes with quadratic losses
In this section we present an alternative strategy developed in [GR09] for empirical Bayes models under the quadratic loss. We begin with a technical result. Let $P_{0},\dots,P_{n}\in{\mathcal{P}}$ be $n+1$ probability distributions, and for $i\in\{0,\dots,n\}$, define the permutation mixture $\mathbb{P}_{-i}$ as
where $S_{A}$ denotes the permutation group over the finite set $A$. In other words, $\mathbb{P}_{-i}$ is the permutation mixture of $\{P_{0},\dots,P_{n}\}\backslash\{P_{i}\}$. The idea of [GR09] leads to the following inequality (though originally stated in a more complicated way):
Lemma 5.4.
For ¯$\overline{\mathbb{P}}=\frac{1}{n+1}\sum_{i=0}^{n}\mathbb{P}_{-i}$ and every $i\in\{0,\dots,n\}$ it holds that
In particular, for every $i,j\in\{0,\dots,n\}$ we have
Unlike Theorem 1.6, the proof of Lemma 5.4 relies on a clever coupling between $\overline{\mathbb{P}}$ and $\mathbb{P}_{-i}$ to reduce $\chi^{2}(\overline{\mathbb{P}}\|\mathbb{P}_{-i})$ to the $\chi^{2}$ divergence between a mixture and a simple distribution (the strategy in Section 2.2). Since ${\sf TV}(P,Q)^{2}\leq H^{2}(P,Q)$, Lemma 5.4 indeed provides a better dependence on ${\mathcal{P}}$ in the $O_{{\mathcal{P}}}(\frac{1}{n})$ upper bound of ${\sf TV}(\mathbb{P}_{n},\mathbb{Q}_{n})^{2}$ in Theorem 1.6. However, the results of Theorem 1.6 and Lemma 5.4 are incomparable in general, because of the product measure $\overline{P}^{\otimes n}$ used in Theorem 1.6 which could be easier to handle in the computation of $\mathbb{E}_{\overline{P}^{\otimes n}}[f^{2}]$.
The upper bound on the Hellinger distance in Lemma 5.4 helps to study empirical Bayes models under the quadratic loss. In the compound setting of empirical Bayes above Corollary 1.7, the risk difference $r^{\mathrm{S}}(\theta)-r^{\mathrm{PI}}(\theta)$ could enjoy a smaller upper bound under a quadratic loss $L(\vartheta,\widehat{\vartheta})=\|\vartheta-\widehat{\vartheta}\|_{2}^{2}$. In this case, it is shown in [GR09] that the optimal simple and permutation invariant decision rules are
respectively. In addition, the risk difference admits a simple form
$r^{\mathrm{S}}(\theta)-r^{\mathrm{PI}}(\theta)=n\mathbb{E}\left[\left(\mathbb{E}\left[\vartheta_{1}|X_{1}\right]-\mathbb{E}\left[\vartheta_{1}|X\right]\right)^{2}\right].$
For a class of models ${\mathcal{P}}$ (such as the Gaussian location model), it is shown in [GR09] that $r^{\mathrm{S}}(\theta)-r^{\mathrm{PI}}(\theta)=O_{{\mathcal{P}}}(1)$, improving over the $O_{{\mathcal{P}}}(\sqrt{n})$ upper bound in Corollary 1.7 for general losses. The following lemma shows that this result could be extended to more general classes ${\mathcal{P}}$ using Lemma 5.4.
Lemma 5.5.
Let $P_{\theta_{1}},\dots,P_{\theta_{n}}\in{\mathcal{P}}$ with ${\mathsf{D}}_{\chi^{2}}({\mathcal{P}})<\infty$, and $|\theta_{i}|\leq M$ for all $i\in[n]$. Then
$r^{\mathrm{S}}(\theta)-r^{\mathrm{PI}}(\theta)=n\mathbb{E}\left[\left(\mathbb{E}\left[\vartheta_{1}|X_{1}\right]-\mathbb{E}\left[\vartheta_{1}|X\right]\right)^{2}\right]\leq 4M^{2}{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})(1+{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})).$

### v3

@@ 5.1 Divergence as a matrix permanent
Compared with Lemma˜5.3, the result of Lemma˜5.5 has an additional $O(\sqrt{\ell})$ factor, due to the use of Theorem˜4.3 in the complex case. However, such an individual bound is crucial to deal with weighted sums $\sum_{\ell=0}^{n}w_{\ell}S_{\ell}$ in applications such as Theorem˜1.5, as witnessed by the next lemma. Recall the definitions of the neighboring mixtures $\mathbb{P}_{n}$ and $\mathbb{P}_{n}^{\prime}$ in Theorem˜1.5. Similar to (28) and (29), let the matrix $A\in\mathbb{R}^{(n-1)\times(n-1)}$ be constructed based on $P_{2},\ldots,P_{n}\in{\mathcal{P}}$, and $S_{\ell}$ be the degree-$\ell$ homegeneous polynomial in its eigenvalues $(\lambda_{2},\ldots,\lambda_{n-1})$. Applying a similar likelihood ratio computation to Lemma˜5.1 leads to the following upper bound.
Lemma 5.6.
Under the above notations, the following inequality holds:
${\sf TV}(\mathbb{P}_{n},\mathbb{P}_{n}^{\prime})^{2}\leq\frac{1}{4}\int\frac{(\mathrm{d}\mathbb{P}_{n}-\mathrm{d}\mathbb{P}_{n}^{\prime})^{2}}{\mathrm{d}\overline{P}^{\otimes n}}\leq\frac{{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})}{n}\sum_{\ell=1}^{n}\ell S_{\ell-1}.$
We defer the proofs of Lemma˜5.6 and Theorem˜1.5 to Section˜C.13 and Section˜C.4, respectively, and present the proof of Lemma˜5.5 relying again on complex normal random variables. Recalling the notation $A=UDU^{\top}$ from Section˜5.1, let $\widetilde{U}\in\mathbb{R}^{n\times(n-1)}$ be the matrix $U$ with its first column (which is $\frac{\bf 1}{\sqrt{n}}$ by Lemma˜5.2) removed, and $\widetilde{D}=\mathrm{diag}(\lambda_{2},\ldots,\lambda_{n})$ with the first eigenvalue $\lambda_{1}=1$ removed. The following lemma establishes an alternative expression of $S_{\ell}$.
@@
The proof of Lemma˜5.7 again follows from the Isserlis’ theorem and several permutation identities in Lemma˜A.1, and is deferred to Section˜C.14. Now we present the proof of Lemma˜5.5.
Proof of Lemma˜5.5.
Since the matrix $U$ is orthogonal with the first column proportional to $\bf 1$}, all columns of $\widetilde{U}$ are orthogonal to $\bf 1$}. Consequently, the vector $\widetilde{P}z=\widetilde{U}\widetilde{D}^{1/2}z$ sums to zero, so (26) in Theorem˜4.3 gives
@@ 6.1 Tightness of upper bounds
In this section we discuss the tightness of the upper bounds in Theorem˜1.2. Recall that the upper bound (4) is quadratic in ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})$ when ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})$ is small, and the upper bound (5) or (6) is a power function of $\mathsf{\Delta}_{H^{2}}({\mathcal{P}})$ when ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})$ is large. The next lemma shows that this is essentially the right behavior of $\chi^{2}(\mathbb{P}_{n}\|\mathbb{Q}_{n})$ for every family $\mathcal{P}$}.
@@
In the setting of Theorem˜1.2, the following lower bound holds:
@@
Combined with Theorem˜1.2, Lemma˜6.1 shows that when the capacity of the family $\mathcal{P}$} is small, we have ${\mathsf{D}}_{H^{2}}({\mathcal{P}})^{2}\lesssim\sup_{n,P_{1},\dots,P_{n}}\chi^{2}(\mathbb{P}_{n}\|\mathbb{Q}_{n})\lesssim{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})^{2}$, which is often tight as ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})\leq{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})\approx{\mathsf{D}}_{H^{2}}({\mathcal{P}})$ in view of the fact that most $f$-divergences are locally $\chi^{2}$-like [PW25, Chapter 7.10]. At the other extreme, when the family $\mathcal{P}$} is rich, we have $\mathsf{\Delta}_{H^{2}}({\mathcal{P}})^{\Omega(1)}-1\leq\sup_{n,P_{1},\dots,P_{n}}\chi^{2}(\mathbb{P}_{n}\|\mathbb{Q}_{n})\leq(e\mathsf{\Delta}_{H^{2}}({\mathcal{P}}))^{{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})}-1$. Therefore, up to a gap in the exponent, the maximum $H^{2}$ singularity $\mathsf{\Delta}_{H^{2}}({\mathcal{P}})$ plays a central role when it is large. The proof of Lemma˜6.1 computes the variance of a linear test function $\frac{1}{n}\sum_{i=1}^{n}f(X_{i})$ under $\mathbb{P}_{n}$ and $\mathbb{Q}_{n}$.
The next natural question is on the tightness of the exponent ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})$ in the upper bound (5). Using the permanent representation of the $\chi^{2}$ divergence in Lemma˜5.1, we can equivalently ask whether the following bound in Section˜5 is tight:
@@
By Lemma˜6.2, the exponent $\mathsf{C}$} in both the permanent upper bound (31) and the $\chi^{2}$ divergence upper bound (5) in Theorem˜1.2 is tight in the worst case, up to multiplicative constants. However, as opposed to Lemma˜6.1, the statement of Lemma˜6.2 is not pointwise in $\mathcal{P}$}; in other words, this does not mean that the upper bounds of Theorem˜1.2 are always tight for a specific class $\mathcal{P}$}. For example, for the Gaussian family (cf. point 1 of Corollary˜1.3) with $\mu>1$, Theorem˜1.2 and Lemma˜6.1 only imply that $\exp(\Omega(\mu^{2}))\leq\sup_{n,P_{1},\dots,P_{n}}\chi^{2}(\mathbb{P}_{n}\|\mathbb{Q}_{n})\leq\exp(O(\mu^{3}))$, still exhibiting a gap on the exponent. An intuitive reason is that, for specific $\mathcal{P}$}, further properties of the matrix $A$ (besides the trace and spectral gap) could potentially be exploited to lead to better results.
6.2 The permutation channel
We also discuss the implications of our results on the capacity of the noisy permutation channel in information theory. Motivated by mobile networks and DNA coding systems, the noisy permutation channel [Mak20] applies a uniformly random permutation to the outputs of a usual communication channel. Mathematically, let ${\mathcal{P}}=({\mathsf{K}}_{x})_{x\in{\mathcal{X}}}$ be a class of conditional distributions ${\mathsf{K}}_{x}:=P_{Z|X=x}$ (i.e., a channel), and $(Z_{1},\dots,Z_{n})$ be the channel output of an input sequence $(X_{1},\dots,X_{n})$. The final output $(Y_{1},\dots,Y_{n})$ of the permutation channel is a uniformly random permutation of $(Z_{1},\dots,Z_{n})$, i.e., $Y_{i}=Z_{\pi(i)}$ for $\pi\sim\text{\rm Unif}(S_{n})$. To understand and design statistically optimal coding schemes for the permutation channel, a key task is to determine the channel capacity
$C_{n}=\max_{p(x^{n})}I(X^{n};Y^{n}),$ (32)
where the maximum is over all possible distributions of $X^{n}$. Using our result that the distribution $P_{Y^{n}|X^{n}}$ is approximately $\left(\frac{1}{n}\sum_{i=1}^{n}{\mathsf{K}}_{X_{i}}\right)^{\otimes n}$, we obtain the following bounds of $C_{n}$.
@@ 6.2 The permutation channel
Let $\delta({\mathcal{P}}):={\mathsf{C}}_{\chi^{2}}({\mathcal{P}})(1+\mathsf{\Delta}_{H^{2}}({\mathcal{P}}))$, the following bounds hold for $C_{n}$ in (32):
$\sup_{\varepsilon>0}\min\left\{\frac{n\varepsilon^{2}}{4},\log M_{\mathrm{H}}({\mathcal{P}}_{n},\varepsilon)\right\}-\frac{\delta({\mathcal{P}})}{2}-\log 2\leq C_{n}\leq\inf_{\varepsilon>0}\left(n\varepsilon^{2}+\log N_{\mathrm{KL}}({\mathcal{P}}_{n},\varepsilon)\right)+\delta({\mathcal{P}}).$
Here ${\mathcal{P}}_{n}:=\frac{1}{n}({\mathcal{P}}+\dots+{\mathcal{P}})$ with the set addition $A+B:=\{a+b:a\in A,b\in B\}$, and
$N_{\mathrm{KL}}({\mathcal{P}},\varepsilon)$ $:=\min\left\{m:\inf_{{\mathcal{P}}_{0}\subseteq{\mathcal{P}}:|{\mathcal{P}}_{0}|=m}\sup_{P\in{\mathcal{P}}}\min_{P_{0}\in{\mathcal{P}}_{0}}{\sf KL}(P\|P_{0})\leq\varepsilon^{2}\right\},$
$M_{\mathrm{H}}({\mathcal{P}},\varepsilon)$ $:=\max\left\{m:\sup_{{\mathcal{P}}_{0}\subseteq{\mathcal{P}}:|{\mathcal{P}}_{0}|=m}\min_{P,P^{\prime}\in{\mathcal{P}}_{0}:P\neq P^{\prime}}H^{2}(P,P^{\prime})\geq\varepsilon^{2}\right\}$
are the covering and packing numbers under the KL divergence and Hellinger distance, respectively.
In the above result, the quantity $\inf_{\varepsilon>0}\left(n\varepsilon^{2}+\log N_{\mathrm{KL}}(\mathrm{conv}({\mathcal{P}}),\varepsilon)\right)$ is an entropic upper bound on the minimax redundancy of ${\mathcal{P}}_{n}^{\otimes n}$ [HO97, YB99], and the quantity $\sup_{\varepsilon>0}\min\{n\varepsilon^{2}/4,\log M_{\mathrm{H}}({\mathcal{P}}_{n},\varepsilon)\}$ is an entropic lower bound for redundancy [HO97]. Therefore, Lemma˜6.3 shows that an i.i.d. approximation of $P_{Y^{n}|X^{n}}$ is quantitatively accurate up to an ${\mathcal{O}}_{\mathcal{P}}(1)$ additive factor. When $\mathcal{P}$} consists of $d$ linearly independent discrete distributions with strictly positive pmfs, both entropic bounds scale as $\frac{d-1}{2}\log n$, so Lemma˜6.3 recovers the capacity bound $C_{n}=\frac{d-1}{2}\log n+O(1)$ in [Mak20, TP23]. Moreover, unlike the method-of-types technique in [TP23] which requires a finite class $\mathcal{P}$} of discrete distributions, Lemma˜6.3 gives meaningful results even if $|{\mathcal{P}}|=\infty$ or $\mathcal{P}$} is a family of continuous distributions.